TO: Tom Scott
FROM: Nancy Hall
RE: Monograph Acquisitions ¿ Monthly report
DATE: July 13, 2004
With the new fiscal all workflow is back to normal. This resulted in
a total of 1318 items catalogued. As well staff processed a total of
2417 items and placed 1471 orders.
- Staff activities. Staff completed work on
their course in LC classification through Mohawk College Distance Learning.
Feedback has been positive about this exercise and I would recommend
the chance be offered to those who did not participate in this round.
Possibilities may exist
for other courses in this vein in the future. Michele Francis was seconded
for 1 week to HR to work on job evaluation.
¿ We continue to be plagued with issues in HNES such as water shut-offs,
power cuts and network problems and printing.
This has necessitated shuffling of schedules and scrambling for work
for the students on occasion, with them being sent home early one day.
The printing issue seems to be resolved with better printers being installed
there. Hopefully the power outage
and network problems will be resolved to result in more reliable service.
With the new version of Peoplesoft, Linda Gamble and myself are
the only two authenticated users in our department. We are endeavouring
to keep up with US and CDN invoices
as work continues on the interface. Hopefully testing can begin sometime
Our students for this summer are O. Alexander, T. Komher, M. Obwaka
and W. Fok. This month
our students worked on the following projects. Transferred from Scott
storage to HNES storage(clean-up project) ¿ 4512; from Scott to Steacie
(Nursing)- 731; from Scott to Bronfman – 611 and a total of 888
items were weeded from Frost . A project was done
to clean-up item types on storage records for a total of 11,600.